OpenStudy (charlotte123):

A line passes through the point(4,-1) and has a slope of 2/3 . If you use this slope to generate another point on the line, what are the coordinates of that point?

OpenStudy (anonymous):

y+1=2/3 ( x-4)

OpenStudy (anonymous):

x=1. y+1 = 2/3(1-4)= -2 y=-3 (1,-3) is another point.

OpenStudy (anonymous):

The first post is the equation of the line.

OpenStudy (anonymous):

The secone is to find another point on the line.

OpenStudy (charlotte123):

-1,3 sorry but thats not a choice

OpenStudy (anonymous):

from (4, -1) you can go "up 2, right 3" to get another point... and keep doing that to get many others...

OpenStudy (anonymous):

or... "down 2, left 3"